OpenStudy (anonymous):

Two marbles are drawn without replacement from a box with 3 white, 2 green, 2 red and 1 blue marble. Find the probability that one marble is white and one marble is blue.

OpenStudy (anonymous):

how many total marbles?

OpenStudy (anonymous):

you are drawing w/o replacement... does the order matter in this case? that is, if you were to select a white marble first and a blue marble second, is that considered the same as drawing a blue forst and a white second? Does the order matter or just the group (1 white and 1 blue)?

OpenStudy (anonymous):

8 total marbles

OpenStudy (anonymous):

and how many will you select?

OpenStudy (anonymous):

the probability of no red marbles _-15/28 probability of at least on red marble_ 1-15/28=13/28

OpenStudy (anonymous):

order does not matter, (i guess). i must select only two, one first then the other. need to know what the probability is of selecting one white and one blue marble

OpenStudy (anonymous):

all you do is multiply the prob of pulling a white one (3/8)then * that by theprob of pullin a blue(1/7)

OpenStudy (anonymous):

right... order doesn't matter. so long as we get 1 white and 1 blue marble, we have the event we want. so we have \(_8C_2 \) total different groups of 2. that's going to be our denominator. we have 3 white marbles, of which we need to choose 1 and 1 blue marble, of which we need to choose 1. so we get this for our probability:\[P \left( 1 \text{ white and }1\text{ blue} \right)=\frac{ \left(\begin{matrix}3 \\ 1\end{matrix}\right) \left(\begin{matrix}1 \\ 1\end{matrix}\right)}{ \left(\begin{matrix}8 \\ 2\end{matrix}\right) }\]

OpenStudy (anonymous):

do you know how to compute this?

OpenStudy (anonymous):

@saue7414 that is incorrect because that is the probability of selecting a white 1st and then a blue. there is another sequence which also satisfies the requirement... blue 1st then a white. the combinations account for this, your method doesn't.

OpenStudy (anonymous):

okay thanks for the correction that would be if order mattered right?

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

ohhhhhhhh LOL

OpenStudy (anonymous):

All i can see from the question is "Find the probability that one marble is white and one marble is blue." Does that say a white marble MUST be chosen first before the blue marble? that I'M not sure of. But with the question worded as such, it probably makes a difference as the outcome. So to err on the side of caution, I'd say that we would have to go with the probability of selecting a white marble first, PRIOR to selecting a blue marble. because if we selected the blue marble first, that would change the outcomes probability

OpenStudy (anonymous):

Am I correct in assuming that probability of drawing a white marble first, then a blue marble, would be changed if a blue marble was drawn first, followed by the probability of drawing a white marble? I guess that the order would change the probability? Right?

OpenStudy (anonymous):

no... you have to look at the group... the group contains 1 white and 1 blue. the order of selection is irrelevant. so long as 1 white and 1 blue are chosen, you have satisfied the requirement. you can do this in one of 2 ways... you can select a white first, followed by a blue or you can select a bleu first, followed by a white. in terms of the probabilities, we have P({W,B}) = P(W 1st and B 2nd) + P(B 1st and W 2nd) =(3/8)*(1/7) + (1/8)*(3/7) = 2*(3/56) = 3/28 this is the same as\[\frac{ \left(\begin{matrix}3 \\ 1\end{matrix}\right)\left(\begin{matrix}1 \\ 1\end{matrix}\right) }{ \left(\begin{matrix}8 \\ 2\end{matrix}\right) }\]

OpenStudy (anonymous):

If what you are saying is correct, and I have no reason to doubt that, wouldn't there be a distinct difference in probability if 1 white marble, (of which there are three in the box) is drawn after 1 blue marble of which there is only one in the box? Wouldn't the order in which they are drawn efect the probability?